# On world-sheet S-matrix of NSR string in static gauge

### v1

@@ Appendix B Global supersymmetry of expanded action
$\delta_{(1)}X^{i}$ $=-i\partial_{+}X^{i}\partial_{-}X^{j}\psi_{R}^{j}\varepsilon_{L}+i\partial_{-}X^{i}\partial_{+}X^{j}\psi_{L}^{j}\varepsilon_{R}+\frac{1}{2}\psi_{R}^{j}\partial_{+}\psi_{R}^{j}\psi_{L}^{i}\varepsilon_{R}-\frac{1}{2}\psi_{L}^{j}\partial_{-}\psi_{L}^{j}\psi_{R}^{i}\varepsilon_{L}\ ,$
$\delta_{(1)}\psi_{L}^{i}$ $=-(\partial_{+}X^{j})^{2}\partial_{-}X^{i}\varepsilon_{R}-i\partial_{-}X^{j}\partial_{+}\psi_{L}^{i}\psi_{R}^{j}\varepsilon_{L}+i\partial_{-}X^{i}\partial_{+}\psi_{L}^{j}\psi_{L}^{j}\varepsilon_{R}\ ,$
$\delta_{(1)}\psi_{R}^{i}$ $=(\partial_{-}X^{j})^{2}\partial_{+}X^{i}\varepsilon_{L}+i\partial_{+}X^{j}\partial_{-}\psi_{R}^{i}\psi_{L}^{j}\varepsilon_{R}-i\partial_{+}X^{i}\partial_{-}\psi_{R}^{j}\psi_{R}^{j}\varepsilon_{L}\ .$ (B.4)

### v4

@@ Appendix B Global supersymmetry of expanded action
$\delta_{(1)}X^{i}$ $=i\partial_{+}X^{i}\partial_{-}X^{j}\psi_{R}^{j}\varepsilon_{L}-i\partial_{-}X^{i}\partial_{+}X^{j}\psi_{L}^{j}\varepsilon_{R}-\frac{1}{2}\psi_{R}^{j}\partial_{+}\psi_{R}^{j}\psi_{L}^{i}\varepsilon_{R}+\frac{1}{2}\psi_{L}^{j}\partial_{-}\psi_{L}^{j}\psi_{R}^{i}\varepsilon_{L}\ ,$
$\delta_{(1)}\psi_{L}^{i}$ $=(\partial_{+}X^{j})^{2}\partial_{-}X^{i}\varepsilon_{R}+i\partial_{-}X^{j}\partial_{+}\psi_{L}^{i}\psi_{R}^{j}\varepsilon_{L}-i\partial_{-}X^{i}\partial_{+}\psi_{L}^{j}\psi_{L}^{j}\varepsilon_{R}-\frac{i}{2}\partial_{-}X^{j}\partial_{+}\psi_{R}^{j}\psi_{L}^{i}\varepsilon_{L}\ ,$
$\delta_{(1)}\psi_{R}^{i}$ $=-(\partial_{-}X^{j})^{2}\partial_{+}X^{i}\varepsilon_{L}-i\partial_{+}X^{j}\partial_{-}\psi_{R}^{i}\psi_{L}^{j}\varepsilon_{R}+i\partial_{+}X^{i}\partial_{-}\psi_{R}^{j}\psi_{R}^{j}\varepsilon_{L}+\frac{i}{2}\partial_{+}X^{j}\partial_{-}\psi_{L}^{j}\psi_{R}^{i}\varepsilon_{R}\ .$ (B.4)